In a double slit experiment with d=5*a, a=40*\lambda, compare (as a ratio) the intensity of the third order interference maximum with that of the zero order maximum.
Answer #1

the positon of the constructive interference on the screen

y = m*lambda*d/a


a = slit width


d = distance of the sreen from the slit


lambda = wave length


the position of the 3 order maximum = y = (3*lambda*5a)/(a)

y = 15*lambda


the intensity of light on the screen other than the central maxima


I = Io*cos^2(pi*a*y/lambda*a)

Io = intensity of the entral maxima

I / Io = cos^2(pi*a*y/lambda*d)


I/Io = cso^2(pi*a*15*lambda/lambda*5a)


I / Io = cos^2(pi*3) = 0.99